# Improved Bounds for the Freiman-Ruzsa Theorem

### v1

@@ 4 Structure from Polynomial Growth
In this section, we will use Fourier analysis to prove Proposition 2.13. We will use several facts about the Fourier transform and Bohr sets, which are collected in Appendix B. This section is very similar to [SRevisited, Sections 9-10].
@@
We then define $S=\{x\in\operatorname{supp}(X):\text{D}_{\text{KL}}(x+Y||X+Y)\leq 1/8\}$. By Markov’s inequality, $\mathbb{P}(X\in S)\geq 1/2$. By Pinsker’s inequality, for any $x\in S$, $\|p_{x+Y}-p_{X+Y}\|_{\ell^{1}}\leq 1/2$. Thus, for any $x_{1},x_{2}\in S$, $\|p_{x_{1}+Y}-p_{x_{2}+Y}\|_{\ell^{1}}\leq 1$. We then compute
@@
$\sum_{x\in G}p_{Y}(x)\omega\gamma(x)=|\widehat{p_{X+Y}}(\gamma)|\geq 1-\delta.$
@@
Remark. The quantity $\sum_{x\in G}\min(p_{x_{1}+Y}(x),p_{x_{2}+Y}(x))$ is also used in a technique called Bernoulli part decomposition, which has been used to prove local central limit theorems. See [gavalakis2024entropy], [mcdonald] for some examples.
@@
By Ruzsa’s covering lemma [Tao_Vu_2006, Lemma 2.14], $A$ can be covered by at most $K\exp(O(d\log(d)))$ translates of SSS-S, and thus at most $K\exp(O(d\log(d)))$ translates of $M\supseteq S-S$. ∎
@@ 5 Barriers to Improved Bounds
    Find a set $S\subseteq\operatorname{supp}(X)$ such that for all $x_{1},x_{2}\in S$, $\|p_{\ell X+x_{1}}-p_{\ell X+x_{2}}\|_{1}$ is very small. Here, $S$ could be considered to be a set of almost-periods for $\ell X$ in a way that more closely resembles traditional arguments in additive combinatorics, such as in [Bourgain_sumsetAPs] or [CrootSisaskAP].
@@
This observation does not, on its own, prohibit the possibility of using almost-periods as part of a proof of Conjecture 1.5. However, we believe that without any new ideas, the methods used in this paper are unlikely to yield a bound better than $f(K)=O(\log(K)\log\log(K))$ for Question 1.2.
We now turn our attention to a more speculative topic: what kinds of results would be required to prove Conjecture 1.5? Before discussing this, let us again discuss the Gowers-Green-Manners-Tao proof of the conjecture in bounded torsion [GGMTTwoTorsion], [GGMTBoundTorsion]. Although the fibring lemma and its immediate consequences are still true for $G=\mathbb{Z}^{d}$, there are two things present in the bounded torsion case that are missing in the case of unbounded torsion. Both issues are barriers to running a “Ruzsa distance decrement” style argument in general.
Barrier 1: The “endgame” step. if none of your operations, such as passing to sums $X_{1}+X_{2}$ or fibers $X_{1}|X_{1}+X_{2}$ provide a Ruzsa distance decrement, what do you do? In the case of bounded torsion, Gowers-Green-Manners-Tao are able to find another operation that ultimately reduces the Ruzsa distance. This is not possible in $\mathbb{Z}^{d}$: a discrete Gaussian has doubling $\Theta(d)$, which may be very large, and the class of $d$-dimensional Gaussians is stable under both of these operations (as well as many other natural operations one may consider). Moreover, Gaussians in $\mathbb{R}^{d}$ minimize the differential entropic doubling constant, and it is likely (though not known, except for when $d=1$ [TSumsetEntropy, Theorem 1.13]) that they are approximate minimizers in $\mathbb{Z}^{d}$ as well. Thus, the “base case” of a Ruzsa distance induction argument may not necessarily occur when $\text{d}[X_{1};X_{2}]=O(1)$, and whatever argument occurs in a base case must be efficient enough to obtain polynomial bounds for Question 1.2.
Barrier 2: A stable characterization of Ruzsa distance extremizers.In [GMTRevisited], Green-Manners-Tao prove that if $\text{d}[X;Y]\leq\epsilon$ and $\epsilon$ is sufficiently small, then there is a subgroup $H$ of $G$ such that $\text{d}[X;U_{H}]$ and $\text{d}[Y;U_{H}]$ are both at most $12\epsilon.$ This serves as the base case of the Ruzsa distance decrement argument proving Marton’s conjecture in groups with bounded torsion.
Tao proved [TSumsetEntropy, Theorem 1.13] that in $\mathbb{Z}$, for random variables with sufficiently high entropy, that the minimal possible entropic doubling constant is approximately $\frac{1}{2}\log(2)$, which matches the behavior of Gaussians. Let us call a $\mathbb{Z}^{d}$-valued random variable “genuinely $d$-dimensional” if it is not concentrated on any hyperplane in $\mathbb{Z}^{d}$. A possible program for proving Conjecture 1.5 would be to carry out the following two steps:
@@
It is then true that $\text{h}(X_{1}+X_{2})\geq\frac{1}{2}\text{h}(X_{1})+\frac{1}{2}\text{h}(X_{2})+\frac{d}{2}\log(2)$, with equality if and only if $X_{1},X_{2}$ are Gaussian [STAM1959101]. This is known as the Shannon-Stam inequality. However, the question of stability of this inequality is subtle. Courtade, Fathi, and Pananjady [CourtadeFathiPananjady] have shown that for all $\epsilon>0$, there exist $\mathbb{R}$-valued random variables $X_{1},X_{2}$ such that $\text{h}(X_{1}+X_{2})\leq\frac{1}{2}\text{h}(X_{1})+\frac{1}{2}\text{h}(X_{2})+\frac{1}{2}\log(2)+\epsilon$ but that for any Gaussian random variable $Z$, $W_{2}^{2}(X_{i},Z)\geq 1/3$ for $i\in\{1,2\}$, where $W_{2}$ is the Wasserstein 2-distance. The counterexamples are (heavy-tailed) Gaussian mixtures, and they are still close to Gaussians in Ruzsa distance, but this demonstrates that there are limitations to the kinds of methods that can be used to prove stability of the Shannon-Stam inequality in general. There are many results that prove stability under strong hypotheses on the random variables $X_{1}$, $X_{2}$, such as log-concavity (also proved in [CourtadeFathiPananjady]), but it seems unlikely that these kinds of exact conditions can be met by the approximate extremizers obtained in a Ruzsa distance decrement argument. For a survey of some stability estimates for the Shannon-Stam inequality, see [MadimanMelbournePeng, Section 2.2.2]. It is worth mentioning that in the discrete setting $G=\mathbb{Z}$, Gavalakis and Kontoyiannis [gavalakis2025conditionsequalitystabilityshannons] have obtained a stablility estimate for the discrete Shannon-Stam inequality, but only when the random variables involved are log-concave.
Instead of the Shannon-Stam inequality, one may consider other characterizations of Gaussian distributions. One particularly attractive one is the Bernstein-Kac theorem: if $X$ and $Y$ are $\mathbb{R}$-valued random variables such that $X+Y$ and XYX-Y are independent, then $X$ and $Y$ are Gaussian. A number of stability estimates [lukacs1977stability], [bernsteinstability] for this inequality are known, quantifying weak dependence using various metrics, but we do not see at present how to discretize these results and relate them to Ruzsa distance.
@@ Appendix A Entropy and Additive Combinatorics
If $X,Y,Z$ are $G$-valued random variables, we have
$\mathbb{H}(Z-Y)-\mathbb{H}(Y)\leq\sum_{z\in G}p_{Z}(z)\text{D}_{\text{KL}}(z-Y||X).$
Moreover, equality holds if X=ZYX=Z-Y.
This is [GMTRevisited, equation A.18].
@@
Statements(i) and (ii) are from [TSumsetEntropy, Theorem 1.10]. Statement (iii) is [GGMTTwoTorsion, Lemma 5.2]. Statement (iv) is [madiman2008entropy, Theorem I]. Statement (v) follows from subtracting
@@
We have seen that $\text{d}[X;Y]\geq 0$. Equality holds if and only if $X,Y$ are both translates of the uniform distribution on a subgroup $H$ of $G$, as proved in [TSumsetEntropy, Theorem 1.11.(i)].
@@
If a random variable $X$ has small doubling, then it cannot be concentrated on a very small set. The following estimate is not explicitly stated in the literature, but it is used as a step in the proof of [GMTRevisited, Proposition 1.2].
@@
Given the close relationship between the estimates in Proposition A.6 and corresponding combinatorial inequalities well-known in additive combinatorics, one may wonder what the advantage is of working in an entropic framework. One answer to this question is that the entropy notions are considerably better behaved under homomorphisms. This was originally proved in [GMTRevisited], and is the driving force behind the solution to Marton’s conjecture in abelian groups with bounded torsion [GGMTTwoTorsion], [GGMTBoundTorsion]. We record the following estimate, which is [GMTRevisited, Proposition 1.4]. A refinement of this result (explicitly calculating the difference between both sides of the inequality) is used in [GGMTTwoTorsion] and [GGMTBoundTorsion], but we do not need this error term here.
@@ Appendix B Fourier Analysis and Bohr Sets
We will now introduce Bohr sets, which are a type of additively structured set ubiquitous in additive combinatorics. Much of the standard material about Bohr sets can be found in [Tao_Vu_2006, Chapter 4].
@@
In view of this example, it would be useful if we could pass from a Bohr set with statistics like those in Proposition B.7 to a convex coset progression with controlled dimension. The following is [SRevisited, Proposition 9.1], and it allows us to do so efficiently.
@@ A Bohr set $B(S,\rho)$ of rank $d$ is regular if the estimate
Not all Bohr sets are regular, but they can always be modified slightly to obtain a regular Bohr set. The following is [Tao_Vu_2006, Lemma 4.25].
@@
The 1% large spectrum of a subset of an abelian group has some additive structure. We record two variants of this fact: one over the entire group $G$, and one localized to a regular Bohr set. The following two are [Tao_Vu_2006, Lemma 4.36] and [SchoenSisask, Lemma 5.3], respectively.
@@ Appendix C Inverse Theorem for $U^{3}$
In this appendix, we record a sketch of a proof of Theorem 1.8. We follow [GTinverse, Section 9] nearly exactly, except we need to replace two applications of the Bogolyubov Lemma. Throughout, $G$ will be an abelian group of size $N$.
We begin with the following result, which can be used as a substitute for [GTinverse, Proposition 9.1].
@@
By [GTinverse, Lemma 8.3], there is a set $S\subseteq\widehat{\widehat{G}}$ such that $A\cap B(S,1/4)=\{0\}$ and $|S|\leq 1+\log_{2}(L)$. Let $\Psi:\widehat{G}\to(\mathbb{S}^{1})^{S}$ be the homomorphism $((s(\xi))_{s\in S}$. We cover $(\mathbb{S}^{1})^{S}$ by $\exp(O(\log(L)\log(k)))$ cubes with side length $1/4k$. There exists one such cube $Q$ such that $|\{(x,y)\in\Gamma^{\prime}:\Psi(y)\in Q\}|$ has size at least $|\Gamma^{\prime}|\exp(-O(\log L\log k))$. Set $\Gamma^{\prime\prime}=\{(x,y)\in\Gamma^{\prime}:\Psi(y)\in Q$.
@@
The remainder of the proof of this proposition proceeds exactly as in [GTinverse, Lemma 9.1]. ∎
We may then follow the remainder of the arguments from [GTinverse, Section 9]. There are two differences. First, the lower bounds appearing on the right hand side of the equations (9.9) onwards will be of the form $\exp(-O(\log(\eta^{-1})^{1+o(1)}))$ instead of $\eta^{O(1)}$. In turn, one must dilate the radius of the Bohr set by factors of $\exp(-O(\log(\eta^{-1})^{1+o(1)}))$ instead of $\eta^{O(1)}$. The second difference is in the use of the local Bogolyubov lemma at the end of Lemma 9.4. This can be replaced by Lemma B.15, which increases the error term a bit more, but it will still be of the form $\exp(-O(\log(\eta^{-1})^{1+o(1)}))$.

### v2

@@ 4 Structure from Polynomial Growth
In this section, we will use Fourier analysis to prove Proposition 2.13. We will use several facts about the Fourier transform and Bohr sets, which are collected in Appendix B. This section is very similar to [27, Sections 9-10].
@@
We then define $S=\{x\in\operatorname{supp}(X):\text{D}_{\text{KL}}(x+Y||X+Y)\leq 1/8\}$. Since the Kullback-Liebler divergence is always nonnegative, we may apply Markov’s inequality and obtain $\mathbb{P}(X\in S)\geq 1/2$. By Pinsker’s inequality, for any $x\in S$, $\|p_{x+Y}-p_{X+Y}\|_{\ell^{1}}\leq 1/2$. Thus, for any $x_{1},x_{2}\in S$, $\|p_{x_{1}+Y}-p_{x_{2}+Y}\|_{\ell^{1}}\leq 1$. We then compute
@@
$\sum_{x\in G}p_{Y}(x)\omega\overline{\gamma(x)}=|\widehat{p_{Y}}(\gamma)|\geq 1-\delta.$
@@
Remark. The quantity $\sum_{x\in G}\min(p_{x_{1}+Y}(x),p_{x_{2}+Y}(x))$ is also used in a technique called Bernoulli part decomposition, which has been used to prove local central limit theorems. See [7], [20] for some examples.
@@
By Ruzsa’s covering lemma [30, Lemma 2.14], $A$ can be covered by at most $K\exp(O(d\log(d)))$ translates of SSS-S, and thus at most $K\exp(O(d\log(d)))$ translates of $M\supseteq S-S$. ∎
@@ 5 Barriers to Improved Bounds
    Find a set $S\subseteq\operatorname{supp}(X)$ such that for all $x_{1},x_{2}\in S$, $\|p_{\ell X+x_{1}}-p_{\ell X+x_{2}}\|_{1}$ is very small. Here, $S$ could be considered to be a set of almost-periods for $\ell X$ in a way that more closely resembles traditional arguments in additive combinatorics, such as in [2] or [4].
@@
This observation does not, on its own, prohibit the possibility of using almost-periods as part of a proof of Conjecture 1.5. However, we believe that without any new ideas, the methods used in this paper are unlikely to yield a bound better than Theorem 1.4, i.e., $f(K)=O(\log(K)\log\log(K))$, for Question 1.2.
We now turn our attention to a more speculative topic: what kinds of results would be required to prove Conjecture 1.5? Before discussing this, let us again discuss the Gowers-Green-Manners-Tao proof of the conjecture in bounded torsion [10], [9]. Although the fibring lemma and its immediate consequences are still true for $G=\mathbb{Z}^{d}$, there are two things present in the bounded torsion case that are missing in the case of unbounded torsion. Both issues are barriers to running a “Ruzsa distance decrement” style argument in general.
Barrier 1: The “endgame” step. if none of your operations, such as passing to sums $X_{1}+X_{2}$ or fibers $X_{1}|X_{1}+X_{2}$ provide a Ruzsa distance decrement, what do you do? In the case of bounded torsion, Gowers-Green-Manners-Tao are able to find another operation that ultimately reduces the Ruzsa distance. This is not possible in $\mathbb{Z}^{d}$: a discrete Gaussian has doubling $\Theta(d)$, which may be very large, and the class of $d$-dimensional Gaussians is stable under both of these operations (as well as many other natural operations one may consider). Moreover, Gaussians in $\mathbb{R}^{d}$ minimize the differential entropic doubling constant, and it is likely (though not known, except for when $d=1$ [31, Theorem 1.13]) that they are approximate minimizers in $\mathbb{Z}^{d}$ as well. Thus, the “base case” of a Ruzsa distance induction argument may not necessarily occur when $\text{d}[X_{1};X_{2}]=O(1)$, and whatever argument occurs in a base case must be efficient enough to obtain polynomial bounds for Question 1.2.
Barrier 2: A stable characterization of Ruzsa distance extremizers.In [11], Green-Manners-Tao prove that if $\text{d}[X;Y]\leq\epsilon$ and $\epsilon$ is sufficiently small, then there is a subgroup $H$ of $G$ such that $\text{d}[X;U_{H}]$ and $\text{d}[Y;U_{H}]$ are both at most $12\epsilon.$ This serves as the base case of the Ruzsa distance decrement argument proving Marton’s conjecture in groups with bounded torsion.
Tao proved [31, Theorem 1.13] that in $\mathbb{Z}$, for random variables with sufficiently high entropy, that the minimal possible entropic doubling constant is approximately $\frac{1}{2}\log(2)$, which matches the behavior of Gaussians. Let us call a $\mathbb{Z}^{d}$-valued random variable “genuinely $d$-dimensional” if it is not concentrated on any hyperplane in $\mathbb{Z}^{d}$. A possible program for proving Conjecture 1.5 would be to carry out the following two steps:
@@
It is then true that $\text{h}(X_{1}+X_{2})\geq\frac{1}{2}\text{h}(X_{1})+\frac{1}{2}\text{h}(X_{2})+\frac{d}{2}\log(2)$, with equality if and only if $X_{1},X_{2}$ are Gaussian [29]. This is known as the Shannon-Stam inequality. However, the question of stability of this inequality is subtle. Courtade, Fathi, and Pananjady [3] have shown that for all $\epsilon>0$, there exist $\mathbb{R}$-valued random variables $X_{1},X_{2}$ such that $\text{h}(X_{1}+X_{2})\leq\frac{1}{2}\text{h}(X_{1})+\frac{1}{2}\text{h}(X_{2})+\frac{1}{2}\log(2)+\epsilon$ but that for any Gaussian random variable $Z$, $W_{2}^{2}(X_{i},Z)\geq 1/3$ for $i\in\{1,2\}$, where $W_{2}$ is the Wasserstein 2-distance. The counterexamples are (heavy-tailed) Gaussian mixtures, and they are still close to Gaussians in Ruzsa distance, but this demonstrates that there are limitations to the kinds of methods that can be used to prove stability of the Shannon-Stam inequality in general. There are many results that prove stability under strong hypotheses on the random variables $X_{1}$, $X_{2}$, such as log-concavity (also proved in [3]), but it seems unlikely that these kinds of exact conditions can be met by the approximate extremizers obtained in a Ruzsa distance decrement argument. For a survey of some stability estimates for the Shannon-Stam inequality, see [16, Section 2.2.2]. It is worth mentioning that in the discrete setting $G=\mathbb{Z}$, Gavalakis and Kontoyiannis [8] have obtained a stablility estimate for the discrete Shannon-Stam inequality, but only when the random variables involved are log-concave.
Instead of the Shannon-Stam inequality, one may consider other characterizations of Gaussian distributions. One particularly attractive one is the Kac-Bernstein theorem: if $X$ and $Y$ are $\mathbb{R}$-valued random variables such that $X+Y$ and XYX-Y are independent, then $X$ and $Y$ are Gaussian. A number of stability estimates [15], [18] for this inequality are known, quantifying weak dependence using various metrics, but we do not see at present how to discretize these results and relate them to Ruzsa distance.
@@ Appendix A Entropy and Additive Combinatorics
If $Y,Z$ are $G$-valued random variables, we have
$\mathbb{H}(Z-Y)-\mathbb{H}(Y)=\sum_{z\in G}p_{Z}(z)\text{D}_{\text{KL}}(z-Y||Z-Y).$
This is the equality case of [11, equation A.18].
@@
Statements(i) and (ii) are from [31, Theorem 1.10]. Statement (iii) is [10, Lemma 5.2]. Statement (iv) is [17, Theorem I]. Statement (v) follows from subtracting
@@
We have seen that $\text{d}[X;Y]\geq 0$. Equality holds if and only if $X,Y$ are both translates of the uniform distribution on a subgroup $H$ of $G$, as proved in [31, Theorem 1.11.(i)].
@@
If a random variable $X$ has small doubling, then it cannot be concentrated on a very small set. The following estimate is not explicitly stated in the literature, but it is used as a step in the proof of [11, Proposition 1.2].
@@
Given the close relationship between the estimates in Proposition A.6 and corresponding combinatorial inequalities well-known in additive combinatorics, one may wonder what the advantage is of working in an entropic framework. One answer to this question is that the entropy notions are considerably better behaved under homomorphisms. This was originally proved in [11], and is the driving force behind the solution to Marton’s conjecture in abelian groups with bounded torsion [10], [9]. We record the following estimate, which is [11, Proposition 1.4]. A refinement of this result (explicitly calculating the difference between both sides of the inequality) is used in [10] and [9], but we do not need this error term here.
@@ Appendix B Fourier Analysis and Bohr Sets
We will now introduce Bohr sets, which are a type of additively structured set ubiquitous in additive combinatorics. Much of the standard material about Bohr sets can be found in [30, Chapter 4].
@@
In view of this example, it would be useful if we could pass from a Bohr set with statistics like those in Proposition B.7 to a convex coset progression with controlled dimension. The following is [27, Proposition 9.1], and it allows us to do so efficiently.
@@ A Bohr set $B(S,\rho)$ of rank $d$ is regular if the estimate
Not all Bohr sets are regular, but they can always be modified slightly to obtain a regular Bohr set. The following is [30, Lemma 4.25].
@@
The 1% large spectrum of a subset of an abelian group has some additive structure. We record two variants of this fact: one over the entire group $G$, and one localized to a regular Bohr set. The following two are [30, Lemma 4.36] and [28, Lemma 5.3], respectively.
@@ Appendix C Inverse Theorem for $U^{3}$
In this appendix, we record a sketch of a proof of Theorem 1.8. We follow [13, Section 9] nearly exactly, except we need to replace two applications of the Bogolyubov Lemma. Throughout, $G$ will be an abelian group of size $N$.
We begin with the following result, which can be used as a substitute for [13, Proposition 9.1].
@@
By [13, Lemma 8.3], there is a set $S\subseteq\widehat{\widehat{G}}$ such that $A\cap B(S,1/4)=\{0\}$ and $|S|\leq 1+\log_{2}(L)$. Let $\Psi:\widehat{G}\to(\mathbb{S}^{1})^{S}$ be the homomorphism $((s(\xi))_{s\in S}$. We cover $(\mathbb{S}^{1})^{S}$ by $\exp(O(\log(L)\log(k)))$ cubes with side length $1/4k$. There exists one such cube $Q$ such that $|\{(x,y)\in\Gamma^{\prime}:\Psi(y)\in Q\}|$ has size at least $|\Gamma^{\prime}|\exp(-O(\log L\log k))$. Set $\Gamma^{\prime\prime}=\{(x,y)\in\Gamma^{\prime}:\Psi(y)\in Q$.
@@
The remainder of the proof of this proposition proceeds exactly as in [13, Lemma 9.1]. ∎
We may then follow the remainder of the arguments from [13, Section 9]. There are two differences. First, the lower bounds appearing on the right hand side of the equations (9.9) onwards will be of the form $\exp(-O(\log(\eta^{-1})^{1+o(1)}))$ instead of $\eta^{O(1)}$. In turn, one must dilate the radius of the Bohr set by factors of $\exp(-O(\log(\eta^{-1})^{1+o(1)}))$ instead of $\eta^{O(1)}$. The second difference is in the use of the local Bogolyubov lemma at the end of Lemma 9.4. This can be replaced by Lemma B.15, which increases the error term a bit more, but it will still be of the form $\exp(-O(\log(\eta^{-1})^{1+o(1)}))$.